# for Generative QA

### v1

@@ B.1 Missing Experimental Details
.      Dataset      Prompt      LLM      Calibrator      $\text{CE}(h;\beta)$      AUAC      OpenBookQA      Ling1S-Top1      Mistral      BB (ours)      $0.285\pm 0.009$      $0.353\pm 0.043$      OpenBookQA      Ling1S-Top1      Mistral      HS-BB (ours)      $0.211\pm 0.015$      $0.37\pm 0.065$      OpenBookQA      Ling1S-Top1      Mistral      S-BB (ours)      $\mathbf{0.193\pm 0.012}$      $0.34\pm 0.062$      OpenBookQA      Ling1S-Top1      Mistral      S-B      $0.477\pm 0.002$      $0.135\pm 0.004$      OpenBookQA      Ling1S-Top1      Mistral      S      $0.32\pm 0.006$      $0.133\pm 0.005$      OpenBookQA      Ling1S-Top1      Mistral      B      $0.471\pm 0.004$      $0.255\pm 0.023$      OpenBookQA      Ling1S-Top1      Mistral      None      $0.441\pm 0.013$      $\mathbf{0.38\pm 0.015}$      OpenBookQA      Ling1S-Top1      Gemma      BB (ours)      $0.316\pm 0.01$      $0.25\pm 0.028$      OpenBookQA      Ling1S-Top1      Gemma      HS-BB (ours)      $0.221\pm 0.016$      $0.261\pm 0.032$      OpenBookQA      Ling1S-Top1      Gemma      S-BB (ours)      $\mathbf{0.215\pm 0.011}$      $0.257\pm 0.082$      OpenBookQA      Ling1S-Top1      Gemma      S-B      $0.415\pm 0.005$      $0.129\pm 0.009$      OpenBookQA      Ling1S-Top1      Gemma      S      $0.421\pm 0.01$      $0.152\pm 0.013$      OpenBookQA      Ling1S-Top1      Gemma      B      $0.483\pm 0.003$      $0.12\pm 0.017$      OpenBookQA      Ling1S-Top1      Gemma      None      $0.434\pm 0.015$      $\mathbf{0.31\pm 0.017}$      OpenBookQA      Verb1S-Top1      Mistral      BB (ours)      $0.288\pm 0.008$      $0.338\pm 0.035$      OpenBookQA      Verb1S-Top1      Mistral      HS-BB (ours)      $0.232\pm 0.012$      $\mathbf{0.344\pm 0.018}$      OpenBookQA      Verb1S-Top1      Mistral      S-BB (ours)      $\mathbf{0.225\pm 0.011}$      $0.316\pm 0.027$      OpenBookQA      Verb1S-Top1      Mistral      S-B      $0.445\pm 0.006$      $0.331\pm 0.011$      OpenBookQA      Verb1S-Top1      Mistral      S      $0.282\pm 0.01$      $0.3\pm 0.009$      OpenBookQA      Verb1S-Top1      Mistral      B      $0.447\pm 0.003$      $0.28\pm 0.02$      OpenBookQA      Verb1S-Top1      Mistral      None      $0.561\pm 0.018$      $0.311\pm 0.021$      OpenBookQA      Verb1S-Top1      Gemma      S-BB (ours)      $\mathbf{0.228\pm 0.011}$      $\mathbf{0.343\pm 0.027}$      OpenBookQA      Verb1S-Top1      Gemma      S-B      $0.410\pm 0.006$      $0.301\pm 0.011$      OpenBookQA      Verb1S-Top1      Gemma      S      $0.298\pm 0.01$      $0.280\pm 0.011$      OpenBookQA      Verb1S-Top1      Gemma      B      $0.447\pm 0.003$      $0.293\pm 0.04$      OpenBookQA      Verb1S-Top1      Gemma      None      $0.561\pm 0.018$      $0.315\pm 0.036$      OpenBookQA      Verb1S-Top1      Gemma      BB (ours)      $0.261\pm 0.008$      $0.332\pm 0.048$      OpenBookQA      Verb1S-Top1      Gemma      HS-BB (ours)      $\mathbf{0.228\pm 0.012}$      $0.321\pm 0.011$      BigBench      Verb1S-Top1      Mistral      BB (ours)      $0.161\pm 0.003$      $0.665\pm 0.031$      BigBench      Verb1S-Top1      Mistral      HS-BB (ours)      $\mathbf{0.132\pm 0.009}$      $0.672\pm 0.079$      BigBench      Verb1S-Top1      Mistral      S-BB (ours)      $0.134\pm 0.004$      $\mathbf{0.679\pm 0.012}$      BigBench      Verb1S-Top1      Mistral      S-B      $0.361\pm 0.006$      $0.513\pm 0.016$      BigBench      Verb1S-Top1      Mistral      S      $0.208\pm 0.004$      $0.522\pm 0.018$      BigBench      Verb1S-Top1      Mistral      B      $0.351\pm 0.002$      $0.592\pm 0.014$      BigBench      Verb1S-Top1      Mistral      None      $0.248\pm 0.005$      $0.671\pm 0.005$      BigBench      Verb1S-Top1      Gemma      BB (ours)      $0.171\pm 0.003$      $\mathbf{0.503\pm 0.019}$      BigBench      Verb1S-Top1      Gemma      HS-BB (ours)      $\mathbf{0.159\pm 0.006}$      $0.502\pm 0.011$      BigBench      Verb1S-Top1      Gemma      S-BB (ours)      $0.215\pm 0.005$      $0.465\pm 0.024$      BigBench      Verb1S-Top1      Gemma      S-B      $0.456\pm 0.001$      $0.289\pm 0.011$      BigBench      Verb1S-Top1      Gemma      S      $0.259\pm 0.004$      $0.292\pm 0.007$      BigBench      Verb1S-Top1      Gemma      B      $0.431\pm 0.002$      $0.441\pm 0.014$      BigBench      Verb1S-Top1      Gemma      None      $0.458\pm 0.01$      $0.499\pm 0.01$
Appendix C KD-tree construction for $\beta$-calibration
@@
We adapt kd-tree [Bentley, 1975] to construct our $\beta$ partitions. We first recall the standard construction of a kd-tree. Let $Z_{i}=(Z_{i1},\ldots,Z_{im},\ldots,Z_{iM}),i=1,\ldots,N$ be the $M$-dimensional dataset to bin and let $H^{k}$ with $k=0$ represent $\{Z_{1},\ldots,Z_{N}\}$. Let $z^{k}_{m}$ denote the median of all the $m$th coordinates of the $Z$s in $H^{k}$. Let $Z[p]$ denote the $p$th coordinate of a $Z$ vector. Set base case value of $k=0$.
@@
In our setting, we form bounded spaces so that our calibrators that contain a $\beta$-binning subroutine can generalize well during test time (outliers that are the closest to the boundary of a space will not be assigned to that space). We take the observations with the smallest and largest order statistic in each coordinate used for pivoting, and use them as bounding values for that coordinate.

### v2

@@ B.1 Missing Experimental Details
. Dataset Prompt LLM Calibrator $\text{CE}(h;\beta)$ AUAC OpenBookQA Ling1S-Top1 Mistral QAB (ours) $0.285\pm 0.009$ $0.353\pm 0.043$ OpenBookQA Ling1S-Top1 Mistral HS-QAB (ours) $0.211\pm 0.015$ $0.37\pm 0.065$ OpenBookQA Ling1S-Top1 Mistral S-QAB (ours) $\mathbf{0.193\pm 0.012}$ $0.34\pm 0.062$ OpenBookQA Ling1S-Top1 Mistral S-B $0.477\pm 0.002$ $0.135\pm 0.004$ OpenBookQA Ling1S-Top1 Mistral S $0.32\pm 0.006$ $0.133\pm 0.005$ OpenBookQA Ling1S-Top1 Mistral B $0.471\pm 0.004$ $0.255\pm 0.023$ OpenBookQA Ling1S-Top1 Mistral None $0.441\pm 0.013$ $\mathbf{0.38\pm 0.015}$ OpenBookQA Ling1S-Top1 Gemma QAB (ours) $0.316\pm 0.01$ $0.25\pm 0.028$ OpenBookQA Ling1S-Top1 Gemma HS-QAB (ours) $0.221\pm 0.016$ $0.261\pm 0.032$ OpenBookQA Ling1S-Top1 Gemma S-QAB (ours) $\mathbf{0.215\pm 0.011}$ $0.257\pm 0.082$ OpenBookQA Ling1S-Top1 Gemma S-B $0.415\pm 0.005$ $0.129\pm 0.009$ OpenBookQA Ling1S-Top1 Gemma S $0.421\pm 0.01$ $0.152\pm 0.013$ OpenBookQA Ling1S-Top1 Gemma B $0.483\pm 0.003$ $0.12\pm 0.017$ OpenBookQA Ling1S-Top1 Gemma None $0.434\pm 0.015$ $\mathbf{0.31\pm 0.017}$ OpenBookQA Verb1S-Top1 Mistral QAB (ours) $0.288\pm 0.008$ $0.338\pm 0.035$ OpenBookQA Verb1S-Top1 Mistral HS-QAB (ours) $0.232\pm 0.012$ $\mathbf{0.344\pm 0.018}$ OpenBookQA Verb1S-Top1 Mistral S-QAB (ours) $\mathbf{0.225\pm 0.011}$ $0.316\pm 0.027$ OpenBookQA Verb1S-Top1 Mistral S-B $0.445\pm 0.006$ $0.331\pm 0.011$ OpenBookQA Verb1S-Top1 Mistral S $0.282\pm 0.01$ $0.3\pm 0.009$ OpenBookQA Verb1S-Top1 Mistral B $0.447\pm 0.003$ $0.28\pm 0.02$ OpenBookQA Verb1S-Top1 Mistral None $0.561\pm 0.018$ $0.311\pm 0.021$ OpenBookQA Verb1S-Top1 Gemma S-QAB (ours) $\mathbf{0.228\pm 0.011}$ $\mathbf{0.343\pm 0.027}$ OpenBookQA Verb1S-Top1 Gemma S-B $0.410\pm 0.006$ $0.301\pm 0.011$ OpenBookQA Verb1S-Top1 Gemma S $0.298\pm 0.01$ $0.280\pm 0.011$ OpenBookQA Verb1S-Top1 Gemma B $0.447\pm 0.003$ $0.293\pm 0.04$ OpenBookQA Verb1S-Top1 Gemma None $0.561\pm 0.018$ $0.315\pm 0.036$ OpenBookQA Verb1S-Top1 Gemma QAB (ours) $0.261\pm 0.008$ $0.332\pm 0.048$ OpenBookQA Verb1S-Top1 Gemma HS-QAB (ours) $\mathbf{0.228\pm 0.012}$ $0.321\pm 0.011$ BigBench Ling1S-Top1 Mistral QAB (ours) $0.195\pm 0.002$ $0.622\pm 0.009$ BigBench Ling1S-Top1 Mistral HS-QAB (ours) $\mathbf{0.138\pm 0.006}$ $\mathbf{0.690\pm 0.031}$ BigBench Ling1S-Top1 Mistral S-QAB (ours) $0.157\pm 0.010$ $0.641\pm 0.02$ BigBench Ling1S-Top1 Mistral S-B $0.387\pm 0.006$ $0.523\pm 0.006$ BigBench Ling1S-Top1 Mistral S $0.224\pm 0.003$ $0.52\pm 0.004$ BigBench Ling1S-Top1 Mistral B $0.379\pm 0.003$ $0.612\pm 0.017$ BigBench Ling1S-Top1 Mistral None $0.245\pm 0.004$ $0.684\pm 0.007$ BigBench Ling1S-Top1 Gemma QAB (ours) $0.198\pm 0.006$ $0.41\pm 0.012$ BigBench Ling1S-Top1 Gemma HS-QAB (ours) $\mathbf{0.142\pm 0.009}$ $\mathbf{0.421\pm 0.021}$ BigBench Ling1S-Top1 Gemma S-QAB (ours) $0.193\pm 0.002$ $0.321\pm 0.052$ BigBench Ling1S-Top1 Gemma S-B $0.326\pm 0.002$ $0.302\pm 0.005$ BigBench Ling1S-Top1 Gemma S $0.180\pm 0.006$ $0.298\pm 0.005$ BigBench Ling1S-Top1 Gemma B $0.468\pm 0.001$ $0.293\pm 0.014$ BigBench Ling1S-Top1 Gemma None $0.427\pm 0.006$ $0.412\pm 0.01$ BigBench Verb1S-Top1 Mistral QAB (ours) $0.161\pm 0.003$ $0.665\pm 0.031$ BigBench Verb1S-Top1 Mistral HS-QAB (ours) $\mathbf{0.132\pm 0.009}$ $0.672\pm 0.079$ BigBench Verb1S-Top1 Mistral S-QAB (ours) $0.134\pm 0.004$ $\mathbf{0.679\pm 0.012}$ BigBench Verb1S-Top1 Mistral S-B $0.361\pm 0.006$ $0.513\pm 0.016$ BigBench Verb1S-Top1 Mistral S $0.208\pm 0.004$ $0.522\pm 0.018$ BigBench Verb1S-Top1 Mistral B $0.351\pm 0.002$ $0.592\pm 0.014$ BigBench Verb1S-Top1 Mistral None $0.248\pm 0.005$ $0.671\pm 0.005$ BigBench Verb1S-Top1 Gemma QAB (ours) $0.171\pm 0.003$ $\mathbf{0.503\pm 0.019}$ BigBench Verb1S-Top1 Gemma HS-QAB (ours) $\mathbf{0.159\pm 0.006}$ $0.502\pm 0.011$ BigBench Verb1S-Top1 Gemma S-QAB (ours) $0.215\pm 0.005$ $0.465\pm 0.024$ BigBench Verb1S-Top1 Gemma S-B $0.456\pm 0.001$ $0.289\pm 0.011$ BigBench Verb1S-Top1 Gemma S $0.259\pm 0.004$ $0.292\pm 0.007$ BigBench Verb1S-Top1 Gemma B $0.431\pm 0.002$ $0.441\pm 0.014$ BigBench Verb1S-Top1 Gemma None $0.458\pm 0.01$ $0.499\pm 0.01$
Appendix C KD-tree construction for QA-calibration
@@
We adapt kd-tree (Bentley, 1975) to construct our $\beta$ partitions. We first recall the standard construction of a kd-tree. Let $Z_{i}=(Z_{i1},\ldots,Z_{im},\ldots,Z_{iM}),i=1,\ldots,N$ be the $M$-dimensional dataset to bin and let $H^{k}$ with $k=0$ represent $\{Z_{1},\ldots,Z_{N}\}$. Let $z^{k}_{m}$ denote the median of all the $m$th coordinates of the $Z$s in $H^{k}$. Let $Z[p]$ denote the $p$th coordinate of a $Z$ vector. Set base case value of $k=0$.
@@
In our setting, we form bounded spaces so that our calibrators that contain a QA binning subroutine can generalize well during test time (outliers that are the closest to the boundary of a space will not be assigned to that space). We take the observations with the smallest and largest order statistic in each coordinate used for pivoting, and use them as bounding values for that coordinate.
@@
Appendix D Generalizing maximum calibration error (MCE) to $\beta$ maximum calibration error (QA MCE)
Guo et al. (2017) defined the maximum calibration error (MCE) as:
$\text{MCE}(h;\beta)=\sup_{r\in\text{range}(h)}\left[\left|\mathbb{E}[Y\mid h(Q,A)=r]-r\right|\right].$
We generalize this definition by incorporating $\beta$:
Definition D.1 (QA Maximum Calibration Error).
The QA maximum calibration error (QA MCE) of $h$ is defined as:
$\beta\text{-MCE}(h;\beta)=\max_{s\in\text{range}(\beta)}\sup_{r\in\text{range}(h)}\left[\left|\mathbb{E}[Y\mid h(Q,A)=r,\beta(Q,A)=s]-r\right|\right].$
[…]
Appendix E Qualitative examples of partition content
Examples of question-and-answer pairs assigned to $\beta$ partitions are presented in Table 8.
Table 8: Qualitative examples of three random partitions and four random question-and-answer pairs from the OpenBookQA dataset, a multiple-choice dataset. The prompt used is Ling1STop1 (Table 5) and answers are generated using Mistral. The partitioning function $\beta$ leverages DistilBERT embeddings and a kd-tree with a maximum depth of 7. The underlying topic for the first two partitions seem clear: “water” and “animal”, and the third partition is likely to have identified the form of the question.
Question LLM Answer If weather is stormy then there is a greater chance of everything getting wet Bill missed high tide, so he had to wait until when to see it again tomorrow Great lakes may have come to be thanks to ice pillars If a flood is occurring, there was most likely great droplets repeating Some animals are aided in finding food sources by aroma Flowers make themselves attractive to hummingbirds with an optimal angle Hummingbirds gather nectar using their bills An example of an adult animal laying eggs is all aside from kittens When a human’s organs stop working and he stops breathing, that person perished If a thing experiences a burning combustion, then it is damaged An exertion on a thing that is going against the thing’s intended direction, when in motion will oppose it When we think of bees, we also think of pollen. This is because bees consume it
Appendix F Extra experiments with varying $\beta$ choices
In addition to using kd-tree as $\beta_{\text{bin}}$ and DistilBERT embedding as $\beta_{\text{emb}}$, we also use k-means as $\beta_{\text{bin}}$ and XLNet (Yang et al., 2019) embeddings as $\beta_{\text{emb}}$ We utilize the [CLS] token embedding from a pre-trained XLNet and conduct a hyperparameter search over the number of centroids, using AUAC as the validation metric. We present the results in Table 9. Similar patterns to those in Table 2 are observed: our methods, S-QAB and HS-QAB, achieve the highest performance for $\text{CE}(h;\beta)$ and AUAC.
[…]
. Dataset Prompt LLM Calibrator $\beta_{\text{emb}}$ $\beta_{\text{bin}}$ $\text{CE}(h;\beta)$ AUAC OpenBookQA Verb1S-Top1 Mistral QAB (ours) DistilBERT kd-tree $0.288\pm 0.008$ $0.338\pm 0.035$ OpenBookQA Verb1S-Top1 Mistral HS-QAB (ours) DistilBERT kd-tree $0.232\pm 0.012$ $\mathbf{0.344\pm 0.018}$ OpenBookQA Verb1S-Top1 Mistral S-QAB (ours) DistilBERT kd-tree $\mathbf{0.225\pm 0.011}$ $0.316\pm 0.027$ OpenBookQA Verb1S-Top1 Mistral S-B DistilBERT kd-tree $0.445\pm 0.006$ $0.331\pm 0.011$ OpenBookQA Verb1S-Top1 Mistral S DistilBERT kd-tree $0.282\pm 0.01$ $0.3\pm 0.009$ OpenBookQA Verb1S-Top1 Mistral B DistilBERT kd-tree $0.447\pm 0.003$ $0.28\pm 0.02$ OpenBookQA Verb1S-Top1 Mistral None DistilBERT kd-tree $0.561\pm 0.018$ $0.311\pm 0.021$ OpenBookQA Verb1S-Top1 Mistral QAB (ours) DistilBERT k-means $0.310\pm 0.009$ $0.334\pm 0.012$ OpenBookQA Verb1S-Top1 Mistral HS-QAB (ours) DistilBERT k-means $\mathbf{0.240\pm 0.013}$ $\mathbf{0.340\pm 0.019}$ OpenBookQA Verb1S-Top1 Mistral S-QAB (ours) DistilBERT k-means $0.255\pm 0.012$ $0.321\pm 0.019$ OpenBookQA Verb1S-Top1 Mistral S-B DistilBERT k-means $0.450\pm 0.007$ $0.331\pm 0.011$ OpenBookQA Verb1S-Top1 Mistral S DistilBERT k-means $0.285\pm 0.011$ $0.3\pm 0.009$ OpenBookQA Verb1S-Top1 Mistral B DistilBERT k-means $0.452\pm 0.004$ $0.28\pm 0.02$ OpenBookQA Verb1S-Top1 Mistral None DistilBERT k-means $0.565\pm 0.018$ $0.311\pm 0.021$ OpenBookQA Verb1S-Top1 Mistral QAB (ours) XLNet kd-tree $0.295\pm 0.010$ $0.337\pm 0.034$ OpenBookQA Verb1S-Top1 Mistral HS-QAB (ours) XLNet kd-tree $0.238\pm 0.012$ $0.341\pm 0.021$ OpenBookQA Verb1S-Top1 Mistral S-QAB (ours) XLNet kd-tree $\mathbf{0.220\pm 0.011}$ $\mathbf{0.347\pm 0.017}$ OpenBookQA Verb1S-Top1 Mistral S-B XLNet kd-tree $0.442\pm 0.008$ $0.331\pm 0.011$ OpenBookQA Verb1S-Top1 Mistral S XLNet kd-tree $0.280\pm 0.012$ $0.3\pm 0.009$ OpenBookQA Verb1S-Top1 Mistral B XLNet kd-tree $0.450\pm 0.004$ $0.28\pm 0.02$ OpenBookQA Verb1S-Top1 Mistral None XLNet kd-tree $0.561\pm 0.018$ $0.311\pm 0.021$ OpenBookQA Verb1S-Top1 Mistral QAB (ours) XLNet k-means $0.312\pm 0.011$ $0.335\pm 0.033$ OpenBookQA Verb1S-Top1 Mistral HS-QAB (ours) XLNet k-means $0.246\pm 0.013$ $0.342\pm 0.019$ OpenBookQA Verb1S-Top1 Mistral S-QAB (ours) XLNet k-means $\mathbf{0.235\pm 0.010}$ $\mathbf{0.350\pm 0.018}$ OpenBookQA Verb1S-Top1 Mistral S-B XLNet k-means $0.448\pm 0.007$ $0.331\pm 0.011$ OpenBookQA Verb1S-Top1 Mistral S XLNet k-means $0.283\pm 0.011$ $0.3\pm 0.009$ OpenBookQA Verb1S-Top1 Mistral B XLNet k-means $0.455\pm 0.005$ $0.28\pm 0.02$ OpenBookQA Verb1S-Top1 Mistral None XLNet k-means $0.565\pm 0.018$ $0.311\pm 0.021$
Appendix G Scatterplot of Recalibrated Confidence Scores over the Confidence Elicitation Baseline
We examine how the recalibrated confidence scores differ for baseline confidence scores obtained using the Ling1STop1 (Figure 3). The confidence scores obtained by our method QA binning have higher sharpness (Brier, 1950; Gneiting et al., 2007) since we are fitting a UMD fitter per partition. The baseline UMD (B), shown in blue, maps an elicited confidence score from Ling1STop1 to either a single value or a small set of nearby values, due to the delta-randomization in UMD. In contrast, our method, QA binning (QAB), shown in grey, applies a partition-specific UMD that adapts to the specific question-and-answer pair. As a result, the scores calibrated by QA binning span a broader range. This is intuitively reasonable, as a prompt-elicited confidence score of 0.1 from a topic like “geography” may require a different recalibration compared to the one from the “politics”.
Figure 3: Scatter plot and regression lines for posthoc-calibrated scores vs confidence scores from Ling1STop1. Note that confidence scores from Ling1STop1 (shown in the the x-axis) are discretized since confidence statement is drawn from an expression list (Table 5). We use the OpenBookQA dataset and Mistral LLM.
Appendix H Reliability analysis of QA partitioned data
In this section, we demonstrate reliability analysis using ECE and reliability plot through an example. We perform an analysis for each of the four $\beta$ partitions (derived using a kd-tree with a maximum depth of two) in Figure 4. We compare the baselines None (from Ling1S-Top1 prompt) and UMD (which is known to best optimize ECE) against our method QA binnning. Partition-wise analysis reveals that the ECE for each partition (representing question-and-answer pairs that are semantically similar) is best achieved by our method, QA binning. However, performing this reliability analysis becomes increasingly challenging and unwieldy as the cardinality of the range of $\beta$ grows.
Figure 4: The reliability plot compares the baselines None , UMD (B), and our method, QA binning (QAB), across each QA partition. Using the OpenBookQA dataset, Ling1S-Top1 prompt, a Mistral LLM, and a kd-tree with a maximum depth of two, four QA partitions are generated, and we conduct an analysis on each partition. UMD (B) and QA binning (QAB), provide confidence scores that are better calibrated at each of the partition compared to None. However, QA binning (QAB) yields better calibrated confidence scores than UMD (B).